PC Solutions sells standard keyboards for $84 and extended keyboards for $105. Last week the store sold three times as many standard keyboards as extended. If total sales were $4998, how many of each type were sold?
Found 2 solutions by Boreal, addingup:
Answer by Boreal(15235)

SK=84
EK=105
3SK=EK
Let EK=x
3x*84+105*x=4998
357x=4998
x=14
14 of the $105 keyboards=$1470
42 of the $84 keyboards=$3528
That adds up to $4998

Answer by addingup(3677)

Standard: S
Extended: E
(you can call them x and y if that's what your teacher wants)
-----------------------------------------------------
S= 84
E= 105
84S + 105E= 4998
-----------------------------------------------------
S= 3E In the equation above, let's use this value for E:
84(3E) + 105E= 4998
252E + 105E= 4998
357E= 4998
E= 14 He sold 14 Extended and:
14*3= 42 Standard
---------------------------------------------------------------
Check:
14*105= 1470
42*84= 3528
_ _ _ _-----
Total...4998 We have the right answer